### v1

# Replacing Softmax Similarity with a Sharpened Angular Similarity: Theory and Practice of Scaling To Billion-Context Attention

Abstract
Softmax Attention has a quadratic time complexity in sequence length, which becomes prohibitive to run at long contexts, even with highly optimized GPU kernels. For example, FlashAttention (an exact, GPU-optimized implementation of Softmax Attention) cannot complete a single forward–backward pass of a multi-head attention layer once the context exceeds $\sim\!4$ million tokens on an NVIDIA GH200 (96 GB). We introduce RACE Attention, a kernel-inspired alternative to Softmax Attention that is linear in sequence length and embedding dimension. RACE Attention replaces the exponential kernel with a sharpened angular (cosine) similarity, and approximates attention outputs via randomized projections and soft Locality-Sensitive Hashing (LSH). Across language modeling, masked language modeling, and text/image classification, RACE Attention matches or outperforms strong baselines while reducing wall-clock time and memory. In a controlled scale test, it processes up to 12 million tokens during a single forward-backward pass on an NVIDIA GH200 GPU and 75 million tokens on an Intel Xeon® Gold 5220R CPU—well beyond the practical limits of the current state-of-the-art attention implementations. RACE Attention thus offers a practical, theoretically grounded mechanism for outrageously long context windows on today’s hardware. We hope that it gets adopted in practice.
1 Introduction
The Transformer [31, 11] is the backbone of modern sequence modeling across language, vision [23], and speech [19]. We have seen remarkable improvements over the past few years in reasoning and understanding capabilities. Most of these are attributed to the increased parameters of the transformers along with the capability to process longer context windows than before. All this progress, however, rests on a computationally expensive primitive: Softmax Attention, whose time scales quadratically with context length. As models and contexts grow—from multi-document reasoning to long-form code, audio, and video—this quadratic barrier increasingly dictates who can train and deploy capable systems. Industrial labs mitigate the cost with large-scale distributed hardware; most practitioners cannot. There is a growing need for attention mechanisms that are accurate, fast, and memory-efficient. To highlight the limits of Softmax Attention: even with FlashAttention [10]—the state-of-the-art GPU implementation—a single forward–backward pass of a multi-head attention layer (1 batch, 4 heads, embedding dimension of 128) remains computationally and memory intensive and cannot process sequences beyond $\sim 4$ million tokens on an NVIDIA GH200 (96 GB). Clearly, to achieve an outrageously long context where the target context size is hundreds of millions of tokens or beyond, fundamental rethinking of attention will be required [2].
Linearized and Low-Rank Approximations to Quadratic Attention: Due to the significance of the problem, a very large body of work attempts to accelerate attention by approximating softmax with linear approximations or clever kernel feature maps [15, 6, 24, 25]. Two notable lines of work in the direction of linearly approximating attention are Linear Attention [15] and Performer [6]. Linear Attention replaces the softmax similarity with a simple positive kernel via a feature map, e.g. $\phi(x)=elu(x)+1$. This lets the attention be re-ordered into associative sums, achieving linear computation. Although such a kernel trick reduces computational complexity, it often degrades accuracy, as clearly demonstrated by our experiments in Section 4.1. Performer takes a different approach and cleverly leverages the classical idea of approximating the exponential of an inner product using Random Fourier Features [26]. However, this strategy comes with its own drawbacks. In particular, the method incurs a time complexity that is quadratic in the embedding dimensionality, which offsets many of the intended computational savings. Furthermore, it is well established [1] that approximations based on Random Fourier Features require high-dimensional representations to achieve satisfactory accuracy. Our experimental results in Section 4.2 reinforce this limitation by showing that these methods exhibit poor scalability in practice.
Another category of work replaces the full $N\times N$ attention matrix with a low-rank surrogate. Some methods learn length-wise projections for keys/values (e.g., Linformer), while others use Nyström landmarks to approximate Softmax Attention matrix with a rank-$k$ decomposition (e.g., Nyströmformer). These approaches reduce the leading cost from $\mathcal{O}(N^{2}d)$ to $\mathcal{O}(Nkd)$, at the cost of choosing (and occasionally increasing) $k$ to maintain accuracy. [32, 34]. Moreover, these methods provide no support for autoregressive tasks. As shown in Section 4.1, our method outperforms Linformer in accuracy despite Linformer having $13\%$ more parameters than the other methods. Beyond the empirical shortcomings, a deeper conceptual issue persists: existing approximation approaches lack a rigorous mathematical framework to characterize the trade-offs between efficiency and accuracy. For example, while Performer provides strong kernel-approximation guarantees, a general framework connecting efficiency knobs (e.g., feature count $m$) to downstream accuracy remains limited, and strong accuracy frequently entails large $m$ in practice. As a result, design decisions often appear ad hoc and fragile, leaving methods vulnerable to instability between tasks and settings. Taken together, these limitations explain why, despite the abundance of approximations, Softmax Attention continues to remain the most widely adopted and reliable formulation.
Sparsity is Complementary: We note that there is also a popular line of work [3, 36, 16, 14] that exploits structural information in natural language, with sparsity in attention being among the most widely studied. These approaches are complementary to our proposal, which focus on making the fundamental attention mechanism itself more efficient and mathematically grounded. In principle, our method can be integrated with structural priors such as sparsity to further improve scalability and accuracy. However, since our objective in this paper is to develop fundamentally efficient attention, we will not discuss this line of structural approaches further, instead we view combining them with our method as an important direction for future work.
Our Finding: Standard attention relies on the well-known softmax function, computing
$O=\mathrm{softmax}\!\left(\tfrac{QK^{\top}}{\sqrt{d}}\right)V\,$, (1)
where the softmax is applied row-wise so that attention weights are nonnegative and sum to one.
In this paper, we propose a surprising alternative to the softmax—namely, a high-degree monomial of an angular kernel based on cosine geometry:
$O=\left(1-\left(\frac{\cos^{-1}\!(QK^{\top})}{\pi}\right)\right)^{\gamma}V$ (2)
Eq. (2) should be read as an informal analogue to Eq. (1), where the angular kernel replaces the exponential. A more precise definition, with explicit cosine normalization and row-wise normalization, is given in Section 3.1. We argue that, for sufficiently large values of $\gamma$, this formulation closely mimics the behavior of softmax and refer to it as Angular Attention in the subsequent sections. Importantly, it admits a linear-time approximation algorithm. In particular, we leverage the connection (Section 2.2) between Repeated Arrays-of-Count Estimators (RACE) [8, 18] and the angular kernel to design our algorithm in Section 3.2. We therefore refer to our proposed method as RACE Attention.
RACE Attention is a drop-in replacement for Softmax Attention. We evaluate it in Transformers on language modeling, masked language modeling, and text/image classification (Section 4.1). By reframing similarity around an angular kernel and using differentiable LSH-based sketches, it provides a principled alternative that supports very long contexts on commodity hardware. The sketching view keeps constant factors small: each query mixes with only a fixed bank of $S=LR$ bucket summaries rather than all $N$ keys. Since we never materialize the full attention matrix, the working set stays compact and activation memory drops, enabling much longer sequences with reduced latency. In contrast, Softmax Attention retains full $Q,K,V,O$ tensors, preventing processing of much longer sequences at comparable speed(Section 4.2). In addition to our novel findings about RACE Attention and rigorous supporting experimental evidence, we provide the following:
  • Long-context scaling: We demonstrate scaling at context lengths far beyond prior attention mechanisms—up to 75 million tokens on a CPU and 12 million tokens on a GPU. To the best of our knowledge, this regime has not been reached by existing attention variants on a single device.
  • Trainable RACE: We make the sketch differentiable by using soft assignments to the hypercube corners (rather than hard hashes), enabling end-to-end learning.
  • CPU/GPU pre-training: We support both causal (autoregressive) and non-causal (bidirectional) pre-training on CPU and GPU. For CPU workloads, we provide a custom OpenMP kernel that computes the causal prefix operations in a single pass using efficient streaming algorithm, enabling linear-time, memory-efficient training.
  • Theory and ablations: We provide approximation guarantees in Section 3.3 inherited from LSH and analyze how sketch parameters—number of hash tables $L$, buckets per table RR—govern variance–accuracy trade-offs. Comprehensive ablations quantify accuracy, speed, and memory as these knobs are varied.
2 Background
2.1 Locality-Sensitive Hashing (LSH)
An LSH family $\mathcal{H}$ for a similarity $\mathrm{Sim}$ makes near pairs collide more often than far pairs. Formally, $\mathcal{H}$ is $(S_{0},cS_{0},p_{1},p_{2})$-sensitive if for all $x,y\in\mathbb{R}^{D}$,
$\begin{cases}\mathrm{Sim}(x,y)\geq S_{0}\;\Rightarrow\;\Pr_{h\sim\mathcal{H}}[h(x)=h(y)]\geq p_{1},\\[2.0pt] \mathrm{Sim}(x,y)\leq cS_{0}\;\Rightarrow\;\Pr_{h\sim\mathcal{H}}[h(x)=h(y)]\leq p_{2},\end{cases}\quad p_{1}>p_{2},\;c<1.$
Such families enable sublinear-time approximate nearest-neighbor data structures. A convenient sufficient condition—satisfied by SimHash and WTA [4, 35, 5]—is that the collision probability is a monotone function of similarity, $\Pr_{h\sim\mathcal{H}}[h(x)=h(y)]=f(\mathrm{Sim}(x,y))$ with $f$ increasing.
Figure 1: This figure demonstrates the difference between the linear complexity of RACE Attention and the quadratic complexity of Softmax Attention mechanism. Specifically, we highlight how the final representation $o_{5}$ is computed under Softmax versus RACE. In Softmax, the entire fifth column of the attention score matrix is required. In contrast, RACE does not require the full matrix; instead, it aggregates statistics within LSH-mapped buckets, utilizing the appropriate collision probability $\alpha$ to compute $o_{5}$.
2.2 RACE Sketch
RACE [8, 9] shows that any similarity expressible as a (non-negative) linear combination of LSH collision kernels can be sketched using ACE-style estimation [18]. It provides an unbiased estimator of kernel-density sums for LSH collision kernels and their powers. In particular, ACE/RACE estimates $\sum_{x\in D}k(x,q)^{p}$ by hashing items into counters and reading the counters addressed by the query; averaging across $L$ independent rows reduces variance.
Lemma 1 (Theorem 1 of [8]).
Given a dataset $D$, an LSH family $H$ with finite range $[1,R]$ and a parameter $p$, construct an LSH function $h(x)\to[1,R^{p}]$ by concatenating $p$ independent hashes from $H$. Let $A$ be an ACE array constructed using $h(x)$. Then for any query $q$,
$\mathbb{E}\!\big[A[h(q)]\big]\;=\;\sum_{x\in D}k(x,q)^{p}$
3 Introducing RACE Attention
3.1 Softmax-Like Similarities that Admit Linear-Time Estimation
Given a sequence of $N$ tokens, a Transformer produces for each position $i$ a query $Q_{i}\in\mathbb{R}^{d}$, and for every position $j$ a key $K_{j}\in\mathbb{R}^{d}$ and a value $V_{j}\in\mathbb{R}^{d}$. The output at position $i$ is a weighted average of the values, where the weight on $V_{j}$ reflects the relevance of token $j$ to token $i$. In the standard formulation [31], relevance is computed via the scaled dot product given by Eq. (1). This choice guarantees two useful properties of the attention weights: (i) non-negativity and (ii) they sum to one, so $O_{i}$ is a convex combination of the values. Equally important, the exponential introduces a strong non-linear mapping from similarity scores to attention weights, amplifying small score differences. This observation suggests a broader view: attention weights can be derived from any normalized highly non-linear (exponential like) similarity function. Let $\mathrm{sim}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ be any non-negative similarity function. We can define normalized similarity attention as
$O_{i}\;=\;\frac{\sum_{j=1}^{N}\mathrm{sim}(Q_{i},K_{j})\,V_{j}}{\sum_{j=1}^{N}\mathrm{sim}(Q_{i},K_{j})}$ (3)
Our quest is for finding non-linear (softmax-like) similarity kernels that admit accurate linear-time estimation, eliminating the quadratic cost of attention in both training and inference. We argue that a good starting point is a well known LSHable [8, 7] angular similarity. It is well behaved and normalized, in particular, it depends only on the angle between the vectors $Q_{i}$ and $K_{j}$ and is invariant to their norms: $\mathrm{sim}(Q_{i},K_{j})=1-\cos^{-1}\!\left(\frac{Q_{i}^{\top}K_{j}}{\lVert Q_{i}\rVert\,\lVert K_{j}\rVert}\right)/\pi$
However, unlike exponential in softmax, the raw angular kernel is relatively flat near high similarity values, reducing its ability to sharply discriminate between nearly aligned vectors. To increase contrast, we propose to exponentiate the angular kernel with a sharpening parameter $\gamma$, which accentuates differences among highly similar pairs. After sharpening the kernel, the similarity function becomes as follows:
$\mathrm{sim}(Q_{i},K_{j})=\left(1-\cos^{-1}\!\left(\frac{Q_{i}^{\top}K_{j}}{\lVert Q_{i}\rVert\,\lVert K_{j}\rVert}\right)/\pi\right)^{\gamma}$ (4)
In Figure 2, we show that for sufficiently large $\gamma$, the angular kernel becomes almost indistinguishable from softmax. This is expected because a high degree monomial like $x^{12}$ behaves similarly to an exponential.
Algorithm 1 RACE Attention (non-causal)
Input: $Q,K,V\in\mathbb{R}^{N\times d}$; number of hash tables $L$; number of hyperplanes $P$; temperature $\beta>0$.
Output: $\widehat{O}\in\mathbb{R}^{N\times d}$.
1:for $\ell=1,\dots,L$ do
2:  Draw $W^{(\ell)}\in\mathbb{R}^{P\times d}$ with rows $w_{p}^{(\ell)}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,I_{d})$.
3:  Define the corner set $\mathcal{V}=\{\pm 1\}^{P}$ ($R=2^{P}$) with $v_{r}\in\{\pm 1\}^{P}$.
4:  Build $\Phi_{Q}^{(\ell)},\Phi_{K}^{(\ell)}\in\mathbb{R}^{N\times R}$ with rows
$\phi^{(\ell)}(x)]_{r}=\frac{\exp\{\beta\,(\tanh(W^{(\ell)}x))^{\top}v_{r}\}}{\sum_{r^{\prime}}\exp\{\beta\,(\tanh(W^{(\ell)}x))^{\top}v_{r^{\prime}}\}},\quad x\in\{Q_{i},K_{j}\}.$
5:  Per-table bucket statistics:
$A^{(\ell)}\;=\;(\Phi_{K}^{(\ell)})^{\top}\mathbf{1}_{N}\in\mathbb{R}^{R},\qquad B^{(\ell)}\;=\;(\Phi_{K}^{(\ell)})^{\top}V\in\mathbb{R}^{R\times d}.$
6:end for
7:Compute average across tables: $\texttt{Num}\;=\;\frac{1}{L}\sum_{\ell=1}^{L}\Phi_{Q}^{(\ell)}B^{(\ell)}$ and $\texttt{Den}\;=\;\frac{1}{L}\sum_{\ell=1}^{L}\Phi_{Q}^{(\ell)}A^{(\ell)}$.
8:Return $\widehat{O}\leftarrow\operatorname{diag}(\texttt{Den})^{-1}\,\texttt{Num}$
In its current form, evaluating the attention with similarity given by Eq. (4) is no different from softmax. It naively still requires all $N^{2}$ query–key interactions. Fortunately, any constant exponentiation of angular kernel, belongs to a family, that admits efficient kernel density estimation using RACE sketches [8]. In particular, we will use LSH-based RACE sketches to approximate the kernel in linear time obtaining an algorithmically efficient alternative to Softmax Attention!
Figure 2: Comparison of Softmax and Angular kernels at different sharpening levels $\gamma$. As $\gamma$ (or non-linearity) increases, Angular transitions from flat similarity scores to a sharper distribution, recovering behavior similar to the exponential of the Softmax.
3.2 The Final Algorithm
At a high level, RACE does not approximate the $N\times N$ score matrix (which would remain quadratic). Instead, it sketches the sufficient statistics needed to compute the attention outputs directly, yielding a linear-time approximation. Fig. 1 illustrates this distinction by focusing on $o_{5}$, the output embedding of token 5 after attention. In Softmax Attention, computing $o_{5}$ requires the entire column of the attention matrix to get the weighted combination of vectors. RACE Attention, in contrast, employs LSH-indexed hash functions to assign the $N$ keys and values (Ks and Vs) into $R$ representative bucket summaries. After this assignment, each query (the $Q$) is hashed into buckets under the same LSH scheme. For example, the output for token 5, $o_{5}$, is obtained by mixing the summaries of tokens assigned to the same bucket as the one mapped by $Q_{5}$ under LSH. The mixing is weighted by soft probability values derived from the well-defined collision probability of the LSH mapping function. Averaging across $L$ independent tables, a standard sketching technique, further reduces variance and stabilizes the approximation. A complete step-by-step expansion is provided in Appendix (fig. 6) and Algorithm 1. Furthermore, we also provide an intuitive visualization of how similarities between tokens look like across buckets in Appendix (fig. 7).
We next formalize the RACE Attention mechanism in Algorithm 1. As described in Section 2.2, one final technical hurdle remains: the RACE algorithm is non-differentiable. We get around the non-differentiability of RACE sketches by replacing discrete bucket assignments with soft probabilities and using standard cross-entropy loss, preserving differentiability for end-to-end training. Algorithm 1 consists of three key stages: (i) Soft bucketization: Each query/key $x\in\mathbb{R}^{d}$ is randomly projected via $W^{(\ell)}$ hyperplanes and softly assigned to $R=2^{P}$ corners with distribution $\phi^{(\ell)}(x)$ (steps 2–4), (ii) Bucket aggregation: For each table $\ell$, we form per-bucket statistics by accumulating key weights and their weighted values, namely the mass vector $A^{(\ell)}\in\mathbb{R}^{R}$ and the value-sum matrix $B^{(\ell)}\in\mathbb{R}^{R\times d}$, so that $A^{(\ell)}[r]$ is the total (soft) mass in bucket $r$ and $B^{(\ell)}[r,:]$ is the corresponding sum of values. (step 5), (iii) Global normalization: The algorithm averages across $L$ tables to form $\texttt{Num}=\tfrac{1}{L}\sum_{\ell}\Phi_{Q}^{(\ell)}B^{(\ell)}$ and $\texttt{Den}=\tfrac{1}{L}\sum_{\ell}\Phi_{Q}^{(\ell)}A^{(\ell)}$, and reconstructs the final outputs as $\widehat{O}=\operatorname{diag}(\texttt{Den})^{-1}\,\texttt{Num}$ (steps 7–8).
Computational Complexity: The per-table runtime of Algorithm 1 can be decomposed according to its main steps: Step 2 (random projections) costs $\mathcal{O}(NdP)$, Step 3 (logits over $R=2^{P}$ corners) costs $\mathcal{O}(NPR)$, and Step 5 (bucket aggregation) costs $\mathcal{O}(NRd)$. The global accumulation in Step 7 adds $\mathcal{O}(NRd)$ per table. Thus, the per-table runtime is $\mathcal{O}(NdP+NPR+NRd)=\mathcal{O}(NRd)$, with memory $\mathcal{O}(NR+Rd)$. Across $L$ tables, this becomes $\mathcal{O}(LNRd)$ time and $\mathcal{O}(L(NR+Rd))$ space. Compared to Softmax Attention’s $\mathcal{O}(N^{2}d)$ time and $\mathcal{O}(Nd)$ space (FlashAttention implementation), RACE is more efficient since $R,L\ll N$ and $R,L\ll d$, even for moderate $N$ and $d$.
3.3 Theoretical Analysis of Algorithm 1
Algorithm 1 is presented in terms of random projections, soft bucketization, and per-bucket aggregation. For analysis it is easier to take a kernel approximation view of attention. Each hash table $\ell=1,\dots,L$ induces a randomized feature map $\phi^{(\ell)}:\mathbb{R}^{d}\to\mathbb{R}^{R}$, where $R=2^{P}$ is the number of hypercube corners, and defines the approximate kernel $\widehat{S}^{(\ell)}_{ij}=\big(\phi^{(\ell)}(Q_{i})\big)^{\top}\phi^{(\ell)}(K_{j}).$ Then, averaging across $L$ independent tables yields $\widehat{S}=\tfrac{1}{L}\sum_{\ell=1}^{L}\widehat{S}^{(\ell)}$. This view places soft RACE Attention in the language of kernel methods: it replaces the angular kernel ($\gamma=P$) in Eq. (4) with the randomized sketch $\widehat{S}$ based on LSH-style features. Since $\phi^{(\ell)}(x)$ is a softmax distribution, the approximate kernel $\widehat{S}$ inherits concentration properties from the underlying random Gaussian projections. This allows us to analyze its deviation from the target angular kernel using standard tools from randomized numerical linear algebra (RandNLA). Our analysis requires the following two mild assumptions on the target kernel $S$:
(A1) Row sums of $S$ are bounded away from zero i.e., $s_{\min}:=\min_{i}(S\mathbf{1})_{i}\geq C_{1}N$ for some constant $C_{1}>0$, which ensures stable normalization in attention.
(A2) Spectral norm of $S$ is bounded i.e., $\|S\|_{2}\leq C_{2}N$, which follows from $S_{ij}\in[0,1]$.
Several comments are necessary to better understand the above structural conditions. Condition (A1) rules out degenerate cases where a query has vanishing similarity with all keys, which would make the row-normalization in attention unstable. In practice this assumption is mild: with learned representations, attention rows rarely collapse to near-zero mass, so requiring $s_{\min}\geq cN$ simply rules out degenerate cases where a query is effectively isolated (assigns negligible weight to almost all keys). Condition (A2) is even less restrictive: since $S_{ij}\in[0,1]$, the worst case is attained by the all-ones matrix $J_{N}$, whose spectral norm is exactly $\|J_{N}\|_{2}=N$. Thus bounding $\|S\|_{2}\leq C_{2}N$ merely rules out pathological growth beyond this trivial maximum, and is always satisfied up to a constant factor. We are now ready to state our main quality-of-approximation result:
Theorem 2.
Let $Q,K,V\in\mathbb{R}^{N\times d}$ be the query, key, and value matrices. For parameters $L$, $P$, and $\beta$, and under conditions (A1) and (A2), the estimator $\widehat{O}$ produced by Algorithm 1 satisfies
$\|\widehat{O}-O\|_{\mathrm{rms}}\;=\;\mathcal{O}\Big(\tfrac{P}{\beta}+\sqrt{\tfrac{\log(N/\delta)}{L}}\Big)\,\|V\|_{F}$
with probability at least $1-\delta$. Here, $O\in\mathbb{R}^{N\times d}$ with the $i^{th}$ row $O_{i}$ is defined using Eqs. (3) and (4) with $\gamma=P$, and $\|\widehat{O}-O\|_{\mathrm{rms}}:=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\|\widehat{O}_{i}-O_{i}\|_{2}^{2}}$ denotes per-token root-mean-square (RMS) error between $O$ and $\widehat{O}$.
The bound decomposes into a bias term $\mathcal{O}(P/\beta)$ and a variance term $\mathcal{O}(\sqrt{\log(N/\delta)/L})$. Larger $\beta$ reduces the bias, while increasing $L$ reduces the variance. The dependence on $P$ arises because powering the angular kernel by $P$ makes collisions sharper, but soft bucketization (finite $\beta$) smooths out these decisions and introduces additional bias. To keep this bias small, $\beta$ should be scaled with $P$. In particular, as $\beta,L\to\infty$, the approximation error vanishes. In fact, taking $L=\Theta(\log N)$ prevents the variance from exploding. Together, this kernel reinterpretation provides a precise RandNLA lens for analyzing RACE Attention, with $L$, $P$, and $\beta$ jointly governing its accuracy-efficiency trade-offs. The proof of Theorem 2, together with all intermediate lemmas, is deferred to Appendix 6 due to space constraints.
Remark (Causal masking): Our experiments in Section 4.1 employ RACE Attention with causal masking, implemented efficiently via OpenMP. See Algorithm 2 in the Appendix for the causal soft RACE Attention algorithm. However, our theoretical analysis above applies only to the non-causal setting. Extending the bias–variance guarantees of Theorem 2 to the masked case remains an open problem, as the cumulative-sum constraint interacts non-trivially with the random feature construction. A rigorous analysis of causal RACE Attention is an important direction for future work.
4 Experiments
To avoid cherry-picking and ensure comparability, we adopt the evaluation suites standard in prior efficient-attention work—Linear Attention, Linformer, and Random Feature Attention (RFA) [15, 32, 24]. Concretely, we include text classification (QNLI [27], SST-2 [28], IMDB [20], and Yahoo [37]) to probe moderate-length discriminative accuracy (as in Linformer); autoregressive language modeling to test token-level modeling (as in RFA) on WikiText-103 [22] and Penn Tree Bank (PTB) corpus [21]; masked language modeling on Tiny Stories dataset [13]; image classification (CIFAR-10 [17] and FashionMNIST [33]) to test expressivity using Vision Transformer [12] architecture (as in Linear Attention); and long-context reasoning via Long Range Arena [29] (e.g., ListOps and Text Retrieval) to stress scaling and accuracy. Together these cover four regimes—bidirectional, autoregressive, long-context, and moderate-context text/image classification. Beyond these standard benchmarks, we additionally report extreme-length scaling experiments to tens of millions of tokens. To the best of our knowledge, this is the first work to report experiments with attention spanning close to 100 million tokens. In this section, we introduce an additional hyperparameter $M$ representing the number of ensembles per head i.e., independent replications of the whole sketching scheme inside each attention head. For clarity, FlashAttention is an exact, fused-kernel implementation of Softmax Attention; we therefore use “FlashAttention” and “Softmax Attention” interchangeably when discussing accuracy and runtime.
(a) RACE - $P=2,L=2,M=1$
(b) RACE - $P=3,L=3,M=1$
(c) RACE - $P=4,L=4,M=1$
(d) RACE - $P=2,L=2,M=1$
(e) RACE - $P=3,L=3,M=1$
(f) RACE - $P=4,L=4,M=1$
Figure 3: A rigorous scaling stress-test across hardware. The top row shows GPU scaling results; the bottom row shows CPU scaling results. We run a single forward-backward pass configured with 1 batch, 4 heads, and an embedding dimension of 128. Linformer and Performer use the same low-rank/feature dimension as in Table 1.
Table 1: IMDB @ N=512 results.
Method P L M Accuracy
RACE 1 2 1 81.3%
RACE 2 2 2 80.6%
RACE 3 3 1 81.3%
RACE 4 4 1 81.3%
Linformer-128 78.2%
FlashAttention 80.0%
Angular ($\gamma$=6) 79.6%
Linear 80.9%
Performer-256 81%
Table 2: Tiny Stories (subset) @ N=512
Method P L M Perplexity
RACE 3 4 1 3.9
RACE 4 4 1 3.3
RACE 5 4 2 2.7
RACE 5 5 1 5.1
Linear 6
Angular ($\gamma$ = 8) 2.9
FlashAttention 3.1
Linformer-128 4.6
Performer-256 7.1
Table 3: ListOps @ N=2000
Method P L M Acc.
RACE 2 2 1 41.9%
RACE 2 3 2 41.0%
RACE 3 3 1 41.3%
RACE 4 3 1 41.2%
Linformer-128 38.9%
FlashAttention 41.4%
Angular ($\gamma$=8) 42.2%
Linear 39.6%
Performer-256 40.2%
Table 4: QNLI @ N=2048
Method K L M Accuracy
RACE 2 2 1 60.7%
RACE 3 3 1 60.7%
RACE 4 4 1 61.1%
RACE 5 5 1 60.4%
Linformer-128 60.6%
Linear 60.7%
FlashAttention 61.1%
Angular ($\gamma$=8) 61.7%
Performer-256 61.0%
Table 5: Text Retrieval @ N=8000
Method P L M Acc.
RACE 2 2 1 80.3%
RACE 2 3 1 80.5%
RACE 3 3 1 80.8%
RACE 4 4 1 80.9%
Linformer-128 76.1%
Linear 80.6%
Performer-256 80.8%
Table 6: SST-2 @ N=1024
Method K L M Accuracy
RACE 2 2 1 76.7%
RACE 4 4 1 79.4%
Linformer-128 75.1%
Linear 78 %
FlashAttention 78.5%
Angular ($\gamma$=8) 77.2%
Performer-256 77.3%
Table 7: PTB @ N=128
Method P L M Test PPL
RACE 2 2 1 54.7
RACE 3 3 1 54.2
RACE 4 4 1 53.4
Angular ($\gamma$=8) 58.8
Angular ($\gamma$=12) 57.6
Linear 73.2
FlashAttention 55.4
Table 8: WikiText-103 @ N=1024
Method P L M Test PPL
RACE 2 2 1 23.9
RACE 2 3 1 23.4
RACE 3 3 1 21.9
RACE 3 4 1 21.5
RACE 4 4 1 20.9
FlashAttention 20.9
Angular ($\gamma$=8) 19
Table 9: Sequential CIFAR-10 @ N=1024
Method P L M Accuracy
RACE 2 2 1 63.7%
RACE 3 3 1 62.5%
RACE 3 5 1 65.7%
RACE 4 5 1 65.9%
Linformer-128 63.7%
FlashAttention 61.44%
Angular ($\gamma$=8) 61.69%
Linear 60%
Performer-256 64.9%
Table 10: Sequential FashionMNIST @ N=784
Method P L M Accuracy
RACE 2 5 1 87.7%
RACE 3 5 1 87.5%
RACE 4 4 1 86.6%
RACE 4 5 1 85.7%
Linformer-128 87.7%
FlashAttention 87.2%
Angular ($\gamma$=8) 86.4%
Linear 85.8%
Performer-256 86.6%
Table 11: Tiny Stories @ N=1024
Method P L M Perplexity
RACE 2 2 1 4.2
RACE 2 3 1 4
RACE 3 3 1 3.2
RACE 4 4 1 2.6
Linear 7
Linformer-128 3.7
FlashAttention 2.7
Angular ($\gamma$ = 8) 2.5
FAVOR+ 10
Table 12: Yahoo @ N=256
Method P L M Accuracy
RACE 2 2 1 66.9%
RACE 3 3 1 66.6%
RACE 4 3 1 66.6%
RACE 4 4 1 67.2%
Linformer-128 64.7%
FlashAttention 67.2%
Angular ($\gamma$=8) 67.0%
Linear 66.9%
Performer-256 64.9%
Baselines: We evaluate RACE against widely used baselines with publicly available implementations: FlashAttention [10], Linear Attention [15], Performer [6], and Linformer [32]. These span exact, kernel-linear, and low-rank approximations. All models are tuned per authors’ guidelines and trained under identical settings.
4.1 Is RACE Attention as Accurate as Tansformers?
We report text-classification accuracy in Tables 1, 4, 6, and 12; long-context (LRA) results—ListOps and Text Retrieval—in Tables 3 and 5; image-classification accuracy on CIFAR-10 in Table 9 and 10; and masked language modeling (MLM) perplexity in Tables 2 and 11. For autoregressive language modeling, RACE matches softmax-level perplexity on WikiText-103 and improves upon it on PTB (Tables 7, 8). These results indicate that RACE preserves accuracy in the overlapping regime while delivering consistent gains on long-context settings.
Unless stated otherwise, all methods use the same Transformer backbone (layers, heads, embedding dimension, dropout) and training budget. We train with identical optimizers, schedulers, and batch sizes; full hyperparameters appear in Table 13. Metrics are reported from the best-validation checkpoint. Despite the extra parameters introduced by Linformer’s lengthwise projections, it does not outperform RACE under matched training conditions. For the long-range task in Table 5, FlashAttention and Angular Attention have quadratic time complexity and are prohibitively slow; we therefore exclude them. All experiments were run on NVIDIA A100 GPUs.
4.2 Can we reach 100 million context window on popular hardware?
In this section, we evaluate how RACE Attention scales across common hardware relative to strong baselines. For RACE, we use sketch parameters $(P,L,M)$ chosen to match FlashAttention’s accuracy/perplexity on the same tasks in Section 4.1. For each method, we measure the wall-clock time for a single forward–backward pass of the multi-head attention layer with 1 batch, 4 heads, and embedding dimension of 128, as a function of sequence length, stress-testing context lengths up to 100 million tokens.
(a) GPU scaling
(b) CPU scaling
(c) $d$ scaling @ N=65536N{=}65536
Figure 4: A rigorous scaling stress-test (including FlashAttention) across hardware. Plots (a)–(b) use logarithmic axes. RACE is evaluated with (P=2,L=2,M=1)(P{=}2,L{=}2,M{=}1) throughout; Linformer and Performer use the same low-rank/feature dimension as in Table 1.
How far can we scale attention on standard Intel Xeon® Gold 5220R CPU?
RACE scales up to 75 million tokens for a single forward-backward pass on CPU. By contrast, FlashAttention becomes prohibitively slow at around $\sim\!2$ million tokens due to the quadratic time scaling in sequence length $N$ (see figs. 3 and 4). It is worth noting that FlashAttention does not run out of memory on the CPU DRAM. However, the time required to complete a single forward-backward pass increases quadratically with $N$. RACE is more than $10000\times$ faster than FlashAttention at context length of $33$ million. RACE finishes comfortably under 10 seconds for a single forward-backward pass on this hardware while FlashAttention takes approximately $10^{5}$ seconds on the same hardware. RACE gets even faster with increasing context length and takes about 100 seconds to perform the same operation with $75$ million tokens. This is expected because RACE is linear and FlashAttention is quadratic in the number of tokens. The experiments also highlight that linear attentions’ approximations are not only inaccurate but also significantly slower and have large memory overheads due to large hidden constants. They run about an order of magnitude slower than RACE Attention and even go out of memory at around 33 million context length.
How far can we scale attention on the most powerful GH200 GPU?
An NVIDIA GH200 has 96GB of memory. Here, we observe a similar trend. RACE scales up to 12 million tokens for a single forward-backward pass, whereas FlashAttention becomes impractical around $\sim\!4$ million tokens (see figs. 3 and 4). At $\sim\!4$ million tokens RACE takes merely 0.1 seconds to finish, while FlashAttention needs about 500 seconds, making RACE about $5000\times$ faster on GPUs. While FlashAttention’s activation memory scales linearly with $N$ and $d$, the GPU’s high-bandwidth memory (HBM) is nevertheless exhausted for sufficiently large $N$. This is because, we must retain $Q,K,V,O$ (and their gradients) of size $\mathcal{O}(BHNd)$ with large constants. Even though the $N\times N$ score matrix is never materialized, this footprint exceeds HBM capacity at large $N$, leading to out-of-memory failures (see fig. 4). Furthermore, RACE even scales better on GPU than the cheap but less accurate linear baselines, and they run out of memory around $\sim\!4$ million tokens. RACE handles about $3.5\times$ longer contexts than FlashAttention.
4.3 Right Algorithm beats Hardware Acceleration!
Figure 5: A rigorous scaling test for algorithmic comparison between FlashAttention on GPU vs. RACE Attention on CPU
[…]
5 Discussion
We introduced RACE Attention, a linear-time, memory-efficient alternative to Softmax Attention that estimates a sharpened angular kernel via RACE sketches. By replacing explicit pairwise scores with bucketed aggregates, compute and activation memory scale linearly with context length $N$, and embedding dimension $d$, with constants governed by the sketch parameters $(P,L,M)$. In our rigorous stress-test, we scale at context lengths up to $\mathbf{75\,million}$ tokens on CPU and $\mathbf{12\,million}$ tokens on GPU (single forward–backward pass), a regime infeasible for Softmax Attention and other highly optimized linear attention mechanisms under comparable settings.
At moderate $N$, optimized GPU Softmax Attention (FlashAttention) yields lower wall-clock time due to high parallelism and effective kernel fusion. As $N$ grows, however, quadratic cost and activation memory dominate, capping sequence length and increasing runtime. In contrast, CPU-based RACE maintains linear scaling, enabling both longer contexts and faster speed in the long-context regime by aggregating values within hash buckets. It is promising to extend RACE Attention for inference only scenarios, potentially eliminating the need for a K cache, and to develop a CUDA kernel, analogous to our OpenMP implementation, that scales well for autoregressive tasks on GPU. We leave these explorations for future work.
Acknowledgments
The work was primarily supported by Rice Ken Kennedy Institute (K2I) Generative AI Cluster Funding. We are grateful to Dr. Aditya Desai and Dr. Zhaozhuo Xu for their valuable insights and thought-provoking discussions.
Appendix
Table 13: Experiment Setup and Hyperparameters
Dataset Task Hyperparameters
[…]
Description:
Unless otherwise stated, we use a linear warmup–decay learning-rate schedule. Let $T$ denote the total number of optimizer updates ($T=\mathrm{epochs}\times\mathrm{len(train\_loader)}$)). The learning rate increases linearly from 0 to the base value over the first $0.01T$ updates, then decays linearly to 0 over the remaining T0.01TT-0.01T updates; the scheduler is stepped once per optimizer update. We use the AdamW optimizer with hyperparameters listed in Table 13. For learning rate $2e^{-4}$, we don’t use linear warmup or a scheduler, we simply train the model with the constant learning rate for certain number of epochs given in Table 13.
6 Proof of Theorem 2
This section provides a complete, self-contained theoretical treatment showing that our RACE Attention closely approximates Angular Attention. We give explicit high-probability bounds for (i) the kernel error, (ii) the attention matrix error with a clean separation of numerator vs. denominator effects, and (iii) the end-to-end output error (Theorem 2). Let’s also reintroduce the notations for the convenience of the reader.
6.1 Setup and assumptions
Data. Sequence length $N$, head (per-head) dimension $d$. Queries/keys are unit vectors:
$Q_{i},K_{j}\in\mathbb{R}^{d}\quad\text{with}\quad\|Q_{i}\|_{2}=\|K_{j}\|_{2}=1,\quad i,j\in\{1,\ldots,N\}.$
Target kernel ($P$-powered angular).
$\kappa(Q_{i},K_{j})\;:=\;\kappa_{\mathrm{ang}}(Q_{i},K_{j})^{P}\;=\;\Big(1-\tfrac{1}{\pi}\cos^{-1}(Q_{i}^{\top}K_{j})\Big)^{P}\in[0,1],\qquad S\in\mathbb{R}^{N\times N}\ \text{with}\ S_{ij}=\kappa(Q_{i},K_{j}).$
Soft RACE features. For each ensemble $\ell=1,\dots,L$:
  • Draw $P$ random hyperplanes $W^{(\ell)}\in\mathbb{R}^{P\times d}$ whose rows $w_{t}^{(\ell)}$ are i.i.d.
  • Corners $\mathcal{V}=\{\pm 1\}^{P}$ (size $R=2^{P}$), with corner vectors $v_{r}\in\{\pm 1\}^{P}$.
  • Logits $s^{(\ell)}(x;r):=[\tanh(W^{(\ell)}x)]^{\top}v_{r}$, temperature $\beta>0$.
  • Define the (probability) feature $\phi^{(\ell)}(x)$ by
    $\phi^{(\ell)}(x)]_{r}\;=\;\frac{\exp\{\beta\,s^{(\ell)}(x;r)\}}{\sum_{r^{\prime}}\exp\{\beta\,s^{(\ell)}(x;r^{\prime})\}}.$
RACE kernel and matrices.
For each ensemble, define the per-table kernel matrix
$\widehat{S}^{(\ell)}_{ij}=(\phi^{(\ell)}(Q_{i}))^{\top}(\phi^{(\ell)}(K_{j})),\qquad\widehat{S}=\frac{1}{L}\sum_{\ell=1}^{L}\widehat{S}^{(\ell)}.$
Let the (single-table) bias matrix be $\tilde{B}:=\mathbb{E}[\widehat{S}^{(\ell)}]-S$.
Assumptions.
For convenience, we restate the two assumptions from Section 3.3
  • (A1) Row sums of $S$ are bounded away from zero i.e., $s_{\min}:=\min_{i}(S\mathbf{1})_{i}\geq C_{1}N$ for some constant $C_{1}>0$, which ensures stable normalization in attention.
  • (A2) Spectral norm of of $S$ is bounded i.e., $\|S\|_{2}\leq C_{2}N$, which follows from $S_{ij}\in[0,1]$.
Notation: We denote $\|\cdot\|_{2}$ as spectral norm for a matrix and Euclidean norm for a vector, $\|\cdot\|_{F}$ for the Frobenius norm of a matrix and for a matrix M, we denote $\|M\|_{\infty}=\max_{i}\sum_{j}|M_{ij}|$.
6.2 Kernel construction with the bias term
We begin by formalizing how a single hash table induces a kernel matrix via the soft RACE features. The next lemma records norm properties that will be used repeatedly.
Lemma 3 (Bounds for a single ensemble).
Let $\Phi_{Q}^{(\ell)}\in\mathbb{R}^{N\times R}$ be the matrix with the $i$-th row $\phi^{(\ell)}(Q_{i})^{\top}$ and $\Phi_{K}^{(\ell)}$ defined analogously. Then:
  1. 1.
    $\widehat{S}^{(\ell)}=\Phi_{Q}^{(\ell)}\big(\Phi_{K}^{(\ell)}\big)^{\top}$.
  2. 2.
    Each row of $\Phi_{Q}^{(\ell)}$ and $\Phi_{K}^{(\ell)}$ is a probability vector; hence $\|\Phi_{Q}^{(\ell)}\|_{F},\ \|\Phi_{K}^{(\ell)}\|_{F}\leq\sqrt{N}$.
  3. 3.
    Consequently $\|\widehat{S}^{(\ell)}\|_{F}\leq N$.
Proof.
Each $\phi^{(\ell)}(x)$ is a softmax over $R=2^{P}$ corners, so entries are nonnegative and sum to $1$. Item (1) is by definition of $\widehat{S}^{(\ell)}_{ij}$. For (2), every row $p$ satisfies $\|p\|_{2}\leq\|p\|_{1}=1$, hence $\|\Phi_{Q}^{(\ell)}\|_{F}^{2}=\sum_{i}\|\phi^{(\ell)}(Q_{i})\|_{2}^{2}\leq N$ (and similarly for $\Phi_{K}^{(\ell)}$). Item (3) follows from $\|AB\|_{F}\leq\|A\|_{F}\|B\|_{F}$. ∎
Having controlled the feature-induced matrix norms, we quantify the zero-mean fluctuation of one ensemble around its expectation and prepare moment bounds needed for matrix concentration. Note that the hash projections $W^{(\ell)}$ (and hence $\widehat{S}^{(\ell)}$) are independent for $\ell=1,\dots,L$.
Lemma 4.
Let $X^{(\ell)}:=\widehat{S}^{(\ell)}-\mathbb{E}[\widehat{S}^{(\ell)}]$ and write
$\Delta:=\widehat{S}-S=\frac{1}{L}\sum_{\ell=1}^{L}X^{(\ell)}+\tilde{B}$.
Then:
  1. 1.
    $\mathbb{E}[X^{(\ell)}]=0$.
  2. 2.
    $\|X^{(\ell)}\|_{2}\leq 2N$.
  3. 3.
    With
    $v:=\max\left\{\left\|\sum_{\ell=1}^{L}\mathbb{E}\!\left[\Big(\tfrac{1}{L}X^{(\ell)}\Big)\Big(\tfrac{1}{L}X^{(\ell)}\Big)^{\!\top}\right]\right\|_{2},\;\left\|\sum_{\ell=1}^{L}\mathbb{E}\!\left[\Big(\tfrac{1}{L}X^{(\ell)}\Big)^{\!\top}\Big(\tfrac{1}{L}X^{(\ell)}\Big)\right]\right\|_{2}\right\},$
    we have $v\leq 4N^{2}/L$.
Proof.
(1) By definition, $X^{(\ell)}=\widehat{S}^{(\ell)}-\mathbb{E}[\widehat{S}^{(\ell)}]$, hence $\mathbb{E}[X^{(\ell)}]=\mathbb{E}[\widehat{S}^{(\ell)}]-\mathbb{E}[\widehat{S}^{(\ell)}]=0$.
(2) By Lemma 3(3) we have $\|\widehat{S}^{(\ell)}\|_{2}\leq\|\widehat{S}^{(\ell)}\|_{F}\leq N$. By convexity of the spectral norm,
$\|\mathbb{E}[\widehat{S}^{(\ell)}]\|_{2}\leq\mathbb{E}\big[\|\widehat{S}^{(\ell)}\|_{2}\big]\leq N.$
Therefore, by the triangle inequality,
$\|X^{(\ell)}\|_{2}\;=\;\|\widehat{S}^{(\ell)}-\mathbb{E}[\widehat{S}^{(\ell)}]\|_{2}\;\leq\;\|\widehat{S}^{(\ell)}\|_{2}+\|\mathbb{E}[\widehat{S}^{(\ell)}]\|_{2}\;\leq\;2N.$
(3) Let $Y^{(\ell)}:=\tfrac{1}{L}X^{(\ell)}$. Then
$\sum_{\ell=1}^{L}\mathbb{E}\big[Y^{(\ell)}(Y^{(\ell)})^{\top}\big]=\frac{1}{L^{2}}\sum_{\ell=1}^{L}\mathbb{E}\big[X^{(\ell)}(X^{(\ell)})^{\top}\big].$
Using subadditivity of $\|\cdot\|_{2}$, Jensen, and $\|AB\|_{2}\leq\|A\|_{2}\|B\|_{2}$,
$\left\|\sum_{\ell=1}^{L}\mathbb{E}\big[Y^{(\ell)}(Y^{(\ell)})^{\top}\big]\right\|_{2}\leq\frac{1}{L^{2}}\sum_{\ell=1}^{L}\left\|\mathbb{E}\big[X^{(\ell)}(X^{(\ell)})^{\top}\big]\right\|_{2}\leq\frac{1}{L^{2}}\sum_{\ell=1}^{L}\mathbb{E}\big[\|X^{(\ell)}\|_{2}^{2}\big]\leq\frac{1}{L^{2}}\sum_{\ell=1}^{L}(2N)^{2}=\frac{4N^{2}}{L}.$
The same bound holds for $\left\|\sum_{\ell=1}^{L}\mathbb{E}\big[(Y^{(\ell)})^{\top}Y^{(\ell)}\big]\right\|_{2}$ by symmetry. Taking the maximum of the two yields $v\leq 4N^{2}/L$. ∎
To convert the moment and uniform bounds above into a high-probability spectral-norm bound, we invoke a standard matrix Bernstein inequality from [30] , stated next for completeness.
Lemma 5 (Matrix Bernstein).
If $Z^{(\ell)}\in\mathbb{R}^{m\times n}$ are independent mean-zero matrices with $\|Z^{(\ell)}\|_{2}\leq H$ and variance proxy $v$, then for any $t>0$,
$\mathbb{P}\!\left(\left\|\sum_{\ell}Z^{(\ell)}\right\|_{2}\geq t\right)\leq(m+n)\,\exp\!\left(-\frac{t^{2}/2}{v+Ht/3}\right).$
Next, applying Lemma 5 with the parameters established in Lemma 4, we obtain the following nonasymptotic deviation bound for the kernel estimator.
Theorem 6 (Kernel deviation with explicit constants).
With probability at least $1-\delta$,
$\|\widehat{S}-S\|_{2}\;\leq\;\|\tilde{B}\|_{2}\;+\;4\,\frac{N}{\sqrt{L}}\sqrt{\log\!\frac{2N}{\delta}}\;+\;\frac{4}{3}\,\frac{N}{L}\,\log\!\frac{2N}{\delta}.$
Proof.
First, rewrite $\widehat{S}-S$ as
$\widehat{S}-S=\frac{1}{L}\sum_{\ell=1}^{L}\widehat{S}^{(\ell)}-S=\frac{1}{L}\sum_{\ell=1}^{L}\Big(\widehat{S}^{(\ell)}-\mathbb{E}[\widehat{S}^{(\ell)}]\Big)\;+\;\Big(\mathbb{E}[\widehat{S}^{(\ell)}]-S\Big)=\frac{1}{L}\sum_{\ell=1}^{L}X^{(\ell)}+\tilde{B}$.
By the triangle inequality,
$\|\widehat{S}-S\|_{2}\;\leq\;\left\|\frac{1}{L}\sum_{\ell=1}^{L}X^{(\ell)}\right\|_{2}\;+\;\|\tilde{B}\|_{2}.$
It remains to upper bound the random term with high probability.
Set $Z^{(\ell)}:=\frac{1}{L}X^{(\ell)}$. Then the $Z^{(\ell)}$ are independent, mean-zero, $N\times N$ random matrices. From Lemma 4(2) we have $\|X^{(\ell)}\|_{2}\leq 2N$. Therefore, $\|Z^{(\ell)}\|_{2}\leq H:=\frac{2N}{L}$. Similarly, Lemma 4(3) gives $v\leq\frac{4N^{2}}{L}$. Applying Lemma 5 with $m=n=N$ yields
$\mathbb{P}\!\left(\left\|\sum_{\ell=1}^{L}Z^{(\ell)}\right\|_{2}\geq t\right)\;\leq\;2N\,\exp\!\left(-\frac{t^{2}}{2\,(v+Ht/3)}\right).$
Let $u:=\log\frac{2N}{\delta}$. To make the RHS $\leq\delta$, it suffices that
$\frac{t^{2}}{2\,(v+Ht/3)}\;\geq\;u\quad\Longleftrightarrow\quad t^{2}-\frac{2uH}{3}\,t-2u\,v\;\geq\;0.$
Choose
$t\;=\;2\sqrt{v\,u}\;+\;\frac{2}{3}\,H\,u.$
Writing $a:=2\sqrt{vu}$ and $b:=\tfrac{2}{3}Hu$ (so $t=a+b$) gives
$t^{2}-\frac{2uH}{3}\,t-2uv=(a+b)^{2}-\frac{2uH}{3}(a+b)-2uv=(4vu-2uv)+\Big(\tfrac{8}{3}-\tfrac{4}{3}\Big)Hu\sqrt{vu}\geq 0.$
Therefore,
$\left\|\sum_{\ell=1}^{L}Z^{(\ell)}\right\|_{2}\;\leq\;2\sqrt{v\,u}\;+\;\frac{2}{3}\,H\,u\quad\text{with probability at least }1-\delta.$
Plugging $v\leq\frac{4N^{2}}{L}$ and $H=\frac{2N}{L}$ yields
$\left\|\sum_{\ell=1}^{L}Z^{(\ell)}\right\|_{2}\;\leq\;2\sqrt{\frac{4N^{2}}{L}\,u}\;+\;\frac{2}{3}\cdot\frac{2N}{L}\,u=4\,\frac{N}{\sqrt{L}}\,\sqrt{u}\;+\;\frac{4}{3}\,\frac{N}{L}\,u.$
Since $\sum_{\ell=1}^{L}Z^{(\ell)}=\frac{1}{L}\sum_{\ell=1}^{L}X^{(\ell)}$, we conclude that
$\left\|\frac{1}{L}\sum_{\ell=1}^{L}X^{(\ell)}\right\|_{2}\;\leq\;4\,\frac{N}{\sqrt{L}}\,\sqrt{\log\!\frac{2N}{\delta}}\;+\;\frac{4}{3}\,\frac{N}{L}\,\log\!\frac{2N}{\delta}\quad\text{with probability }\geq 1-\delta,$
and therefore
$\|\widehat{S}-S\|_{2}\;\leq\;\|\tilde{B}\|_{2}\;+\;4\,\frac{N}{\sqrt{L}}\,\sqrt{\log\!\frac{2N}{\delta}}\;+\;\frac{4}{3}\,\frac{N}{L}\,\log\!\frac{2N}{\delta},$
as claimed. ∎
The deviation bound decomposes into a variance term and a (deterministic) bias term $\tilde{B}$. We now bound $\tilde{B}$ explicitly as a function of $\beta$ and $P$.
Lemma 7 (Bias to the $P$-powered angular kernel: explicit bound).
Fix $P\geq 1$ and $\beta>0$. With $S$ and $\tilde{B}$ as above, let $c:=2\tanh(1)$ and $C_{1}:=\frac{2}{\sqrt{2\pi}}e^{-1/2}$. Then
$\|\tilde{B}\|_{2}\;\leq\;\frac{4}{\sqrt{2\pi}}\;\frac{NP}{\beta}\;+\;\underbrace{\Big(\tfrac{4}{\sqrt{2\pi}}e^{-1/2}\Big)}_{=\,2C_{1}}\,NP\,e^{-c\beta}.$
Proof.
Let us denote the inner product similarity by $\rho:=Q_{i}^{\top}K_{j}$, and recall that the angular kernel is $\kappa_{\mathrm{ang}}(Q_{i},K_{j}):=1-\frac{1}{\pi}\cos^{-1}(\rho)$ From standard LSH theory, for $P$ i.i.d. Gaussian hyperplanes $W^{(\ell)}\in\mathbb{R}^{P\times d}$, the probability that all $P$ bits match is exactly:
$\mathbb{P}\left(h_{P}(Q_{i})=h_{P}(K_{j})\right)=\kappa_{\mathrm{ang}}(Q_{i},K_{j})^{P}=S_{ij}.$
Now define the softmax sketch feature for the hash table $\ell$:
$s^{(\ell)}(x;r):=\tanh(W^{(\ell)}x)^{\top}v_{r},\qquad[\phi^{(\ell)}(x)]_{r}:=\frac{e^{\beta s^{(\ell)}(x;r)}}{\sum_{r^{\prime}}e^{\beta s^{(\ell)}(x;r^{\prime})}},$
where $v_{r}\in\{\pm 1\}^{P}$ denotes the binary corner vectors of length $P$, and $R=2^{P}$. Let $\widehat{S}^{(\ell)}\in\mathbb{R}^{N\times N}$ be the kernel matrix for a single hash table:
$\mathbb{E}(\widehat{S}^{(\ell)}_{ij}):=\mathbb{E}\left[\left(\phi^{(\ell)}(Q_{i})\right)^{\top}\left(\phi^{(\ell)}(K_{j})\right)\right],\qquad S_{ij}:=\kappa_{\mathrm{ang}}(Q_{i},K_{j})^{P},$
and recall the bias matrix is $\tilde{B}:=\mathbb{E}[\widehat{S}^{(\ell)}]-S$. Our goal is to bound $\|\tilde{B}\|_{2}$. To do this, fix any pair $(i,j)$ and note:
$|\mathbb{E}(\widehat{S}^{(\ell)}_{ij})-S_{ij}|=\left|\mathbb{E}\left[\left(\phi^{(\ell)}(Q_{i})\right)^{\top}\left(\phi^{(\ell)}(K_{j})\right)\right]-\mathbb{P}[h_{P}(Q_{i})=h_{P}(K_{j})]\right|.$
Next, let $r^{\star}(x)=\arg\max_{r}s^{(\ell)}(x;r)=\operatorname{sign}(u(x))$ denote the maximizing corner for $x$, where $u_{t}(x)=\tanh(w_{t}^{\top}x)$. The function $s^{(\ell)}(x;r)=\sum_{t=1}^{P}u_{t}(x)\,r_{t}$ is a linear form over the binary corners $r\in\{\pm 1\}^{P}$. To evaluate the normalization term in the softmax, note that the exponentials factorize across coordinates because $r_{t}$ appears only in the term $u_{t}(x)r_{t}$. Hence,
$\sum_{r\in\{\pm 1\}^{P}}e^{\beta s^{(\ell)}(x;r)}$ $=\sum_{r_{1}=\pm 1}\!\cdots\!\sum_{r_{P}=\pm 1}\left(e^{\beta\sum_{t}u_{t}(x)r_{t}}\right)=\sum_{r_{1}=\pm 1}\!\cdots\!\sum_{r_{P}=\pm 1}\prod_{t=1}^{P}e^{\beta u_{t}(x)r_{t}}$
$=\prod_{t=1}^{P}\left(\sum_{r_{t}=\pm 1}e^{\beta u_{t}(x)r_{t}}\right)=\prod_{t=1}^{P}\!\big(e^{\beta u_{t}(x)}+e^{-\beta u_{t}(x)}\big)=\prod_{t=1}^{P}2\cosh(\beta|u_{t}(x)|),$
where the last equality uses the evenness of the hyperbolic cosine, $\cosh(z)=\cosh(|z|)$. Therefore, the softmax probability assigned to the dominant corner $r^{\star}(x)$ can be written in closed form as
$\phi^{(\ell)}(x)]_{r^{\star}(x)}=\frac{e^{\beta\sum_{t}|u_{t}(x)|}}{\prod_{t}2\cosh(\beta|u_{t}(x)|)}=\prod_{t=1}^{P}\frac{e^{\beta|u_{t}(x)|}}{e^{\beta|u_{t}(x)|}+e^{-\beta|u_{t}(x)|}}=\prod_{t=1}^{P}\sigma\!\big(2\beta|u_{t}(x)|\big),$
where $\sigma(z)=1/(1+e^{-z})$ denotes the logistic function. This factorization reveals that each bit contributes independently to the model’s confidence in selecting its sign, and the total probability mass on $r^{\star}(x)$ is the product of these per-bit probabilities.
To bound the total probability mass outside the dominant corner, we use the inequality $1-\prod_{t}(1-a_{t})\leq\sum_{t}a_{t}$ for $a_{t}\in[0,1]$, which we apply to $a_{t}=1-\sigma(2\beta|u_{t}(x)|)$. Hence,
$1-[\phi^{(\ell)}(x)]_{r^{\star}(x)}$ $=1-\prod_{t=1}^{P}\sigma(2\beta|u_{t}(x)|)\leq\sum_{t=1}^{P}\big(1-\sigma(2\beta|u_{t}(x)|)\big)\leq\sum_{t=1}^{P}e^{-2\beta|u_{t}(x)|},$ (5)
where the final inequality follows from the standard logistic bound $1-\sigma(z)\leq e^{-z}$ for all $z\geq 0$. Intuitively, this means that the total softmax probability mass outside the most likely corner decays exponentially with the scaled activation strength $\beta|u_{t}(x)|$ along each coordinate. Now, for each bit, $u_{t}(x)=\tanh(w_{t}^{\top}x)$ with $w_{t}^{\top}x\sim\mathcal{N}(0,1)$. Let $Z:=w_{t}^{\top}x$. Then
$\mathbb{E}[e^{-2\beta|u_{t}(x)|}]=\mathbb{E}[e^{-2\beta|\tanh(Z)|}].$
We split into two regions. (1) On $|Z|\leq 1$, we use $|\tanh z|\geq\frac{|z|}{2}$, so $e^{-2\beta|\tanh(Z)|}\leq e^{-\beta|Z|}$. Hence
$\mathbb{E}[e^{-2\beta|\tanh(Z)|}\mathbf{1}_{|Z|\leq 1}]\leq\frac{2}{\sqrt{2\pi}}\int_{0}^{1}e^{-\beta z}e^{-z^{2}/2}\,dz\leq\frac{2}{\sqrt{2\pi}}\cdot\frac{1}{\beta}.$
(2) On $|Z|>1$, we use $\tanh z\geq\tanh(1)$, so $e^{-2\beta|\tanh(Z)|}\leq e^{-2\beta\tanh(1)}$. Thus
$\mathbb{E}[e^{-2\beta|\tanh(Z)|}\mathbf{1}_{|Z|>1}]\leq e^{-2\beta\tanh(1)}\mathbb{P}(|Z|>1)=2e^{-2\beta\tanh(1)}\,\mathbb{P}(Z>1)\leq e^{-2\beta\tanh(1)}\,\sqrt{\frac{2}{\pi}}e^{-1/2}\leq e^{-c\beta}\,$,
where the second last bound is due to Mill’s inequality and the last inequality follows from the fact that $\sqrt{\frac{2}{\pi}}e^{-1/2}\leq 1$. Here $c=2\tanh(1)$. Combining the two expectations we get,
$\mathbb{E}[e^{-2\beta|u_{t}(x)|}]\leq\frac{2}{\sqrt{2\pi}\,\beta}+e^{-c\beta},\quad c=2\tanh(1).$
Substituting back into Eq. (5), we obtain
$\mathbb{E}\big[1-[\phi^{(\ell)}(x)]_{r^{\star}(x)}\big]\;\leq\;\frac{2P}{\sqrt{2\pi}\,\beta}\;+\;\mathcal{O}(Pe^{-c\beta}).$
Next, let $p:=\phi^{(\ell)}(Q_{i})$ and $q:=\phi^{(\ell)}(K_{j})$ denote the softmax feature vectors for $Q_{i}$ and $K_{j}$, and let $a:=r^{\star}(Q_{i})$, $b:=r^{\star}(K_{j})$ be their respective dominant corners. By the deterministic inequality
$|p^{\top}q-\mathbf{1}\{a=b\}|\;\leq\;(1-p_{a})+(1-q_{b}),$
valid for any probability vectors $p,q$ and indices $a,b$, we have
$|\mathbb{E}[\widehat{S}^{(\ell)}_{ij}]-S_{ij}|\;=\;\big|\mathbb{E}[p^{\top}q]-\mathbb{P}[a=b]\big|\;\leq\;\mathbb{E}[1-p_{a}]+\mathbb{E}[1-q_{b}].$
Using the bound on the expected softmax tail probability for both $Q_{i}$ and $K_{j}$ then gives
$|\mathbb{E}[\widehat{S}^{(\ell)}_{ij}]-S_{ij}|\;\leq\;2\left(\frac{2P}{\sqrt{2\pi}\,\beta}+\mathcal{O}(Pe^{-c\beta})\right)\;=\;\frac{4P}{\sqrt{2\pi}\,\beta}+\mathcal{O}(Pe^{-c\beta}).$
Therefore,
$\|\tilde{B}\|_{2}\leq\|\tilde{B}\|_{F}\leq N\sup_{i,j}|\tilde{B}_{ij}|\;\leq\;N\left(\frac{4P}{\sqrt{2\pi}\,\beta}+\mathcal{O}(Pe^{-c\beta})\right).$
This proves the claim. ∎
We now propagate the kernel-level error into the attention matrix. This requires controlling the normalization (row sums) and its inverse, which we address next.
6.3 From kernels to attention (numerator vs. denominator)
Let $\widehat{S}=S+\Delta$, $D=\operatorname{diag}(S\mathbf{1})$, and $\widehat{D}=\operatorname{diag}(\widehat{S}\mathbf{1})=D+E$. Define the attention matrices
$A:=D^{-1}S,\qquad\widehat{A}:=\widehat{D}^{-1}\widehat{S}$.
The following lemma ties the row-sum perturbation $E$ to $\Delta$ and gives a simple invertibility condition for $\widehat{D}$.
Lemma 8 (Row-sum and inverse diagonal control).
Recall that $s_{\min}=\min_{i}D_{ii}>0$. Then
  1. 1.
    $\|E\|_{2}\leq\|\Delta\|_{\infty}\leq\sqrt{N}\,\|\Delta\|_{2}$.
  2. 2.
    If $\|E\|_{2}\leq s_{\min}/2$, then $\|\widehat{D}^{-1}\|_{2}\leq 2/s_{\min}$.
Proof.
(1) Row-sum control. Since $\widehat{S}=S+\Delta$ and $\widehat{D}=\operatorname{diag}(\widehat{S}\mathbf{1})$, we rewrite
$E\;:=\;\widehat{D}-D\;=\;\operatorname{diag}\big((\widehat{S}-S)\mathbf{1}\big)\;=\;\operatorname{diag}(\Delta\mathbf{1}).$
Hence each diagonal entry is $E_{ii}=(\Delta\mathbf{1})_{i}=\sum_{j=1}^{N}\Delta_{ij}$, so
$\|E\|_{2}\;=\;\max_{i}|E_{ii}|\;=\;\max_{i}\big|(\Delta\mathbf{1})_{i}\big|\;\leq\;\max_{i}\sum_{j=1}^{N}|\Delta_{ij}|\;=\;\|\Delta\|_{\infty}.$
For the second inequality, by Cauchy–Schwarz on each row $i$,
$\sum_{j=1}^{N}|\Delta_{ij}|\;\leq\;\sqrt{N}\,\Big(\sum_{j=1}^{N}\Delta_{ij}^{2}\Big)^{1/2}\;=\;\sqrt{N}\,\|\Delta_{i,\cdot}\|_{2}.$
Moreover,
$\max_{i}\|\Delta_{i,\cdot}\|_{2}=\max_{i}\|\Delta^{\top}e_{i}\|_{2}\;\leq\;\|\Delta^{\top}\|_{2}\|e_{i}\|_{2}\;=\;\|\Delta\|_{2}.$
Taking the maximum over $i$ yields
$\|\Delta\|_{\infty}\;=\;\max_{i}\sum_{j}|\Delta_{ij}|\;\leq\;\sqrt{N}\,\max_{i}\|\Delta_{i,\cdot}\|_{2}\;\leq\;\sqrt{N}\,\|\Delta\|_{2}.$
(2) Inverse diagonal control. Because $\widehat{D}=D+E$ is diagonal, its smallest diagonal entry satisfies
$\min_{i}\widehat{D}_{ii}\;=\;\min_{i}(D_{ii}+E_{ii})\;\geq\;\min_{i}D_{ii}-\max_{i}|E_{ii}|\;=\;s_{\min}-\|E\|_{2}.$
If $\|E\|_{2}\leq s_{\min}/2$, then $\min_{i}\widehat{D}_{ii}\geq s_{\min}/2>0$, so $\widehat{D}$ is invertible and
$\|\widehat{D}^{-1}\|_{2}\;=\;\max_{i}\frac{1}{\widehat{D}_{ii}}\;\leq\;\frac{1}{s_{\min}-\|E\|_{2}}\;\leq\;\frac{1}{s_{\min}-s_{\min}/2}\;=\;\frac{2}{s_{\min}}.$
Assumption (A1) ensures $D$ has diagonals of order $N$, but we must still control $E$. The next lemma shows that the condition $\|E\|_{2}\leq s_{\min}/2$ holds with high probability once $L$ is moderately large.
Lemma 9 (Concentration bound for $E$).
Under assumption (A1), with probability at least $1-\delta$,
$\|E\|_{2}\;\leq\;\tfrac{1}{2}\,s_{\min}$
provided that
$L\;\geq\;\frac{2}{C_{1}^{2}}\,\log\!\frac{2N^{2}}{\delta}.$
Proof.
Recall $E=\widehat{D}-D=\operatorname{diag}(\Delta\mathbf{1})$ with $\Delta=\widehat{S}-S$. Hence
$\|E\|_{2}\;=\;\max_{i}|(\Delta\mathbf{1})_{i}|=\max_{i}\Big|\sum_{j=1}^{N}(\widehat{S}_{ij}-S_{ij})\Big|.$
Each entry $\widehat{S}_{ij}$ is the average of $L$ i.i.d. bounded random variables $\widehat{S}^{(\ell)}_{ij}\in[0,1]$ with mean $S_{ij}$. By Hoeffding’s inequality,
$\Pr\!\big(|\widehat{S}_{ij}-S_{ij}|>\epsilon\big)\;\leq\;2\exp(-2L\epsilon^{2}).$
A union bound over all $N^{2}$ pairs $(i,j)$ gives
$\Pr\!\Big(\max_{i,j}|\widehat{S}_{ij}-S_{ij}|>\epsilon\Big)\;\leq\;2N^{2}\exp(-2L\epsilon^{2}).$
Thus with probability at least $1-\delta$,
$\max_{i,j}|\widehat{S}_{ij}-S_{ij}|\;\leq\;\sqrt{\tfrac{1}{2L}\,\log\!\tfrac{2N^{2}}{\delta}}.$
For any row $i$,
$\Big|\sum_{j=1}^{N}(\widehat{S}_{ij}-S_{ij})\Big|\;\leq\;N\,\max_{j}|\widehat{S}_{ij}-S_{ij}|,$
so with probability $\geq 1-\delta$,
$\|E\|_{2}\;\leq\;N\sqrt{\tfrac{1}{2L}\,\log\!\tfrac{2N^{2}}{\delta}}.$
By (A1), $s_{\min}\geq C_{1}N$. Therefore $\|E\|_{2}\leq\tfrac{1}{2}s_{\min}$ whenever
$N\sqrt{\tfrac{1}{2L}\,\log\!\tfrac{2N^{2}}{\delta}}\;\leq\;\tfrac{1}{2}C_{1}N,$
which simplifies to the claimed condition $L\geq\tfrac{2}{C_{1}^{2}}\,\log\!\tfrac{2N^{2}}{\delta}$. ∎
Now, with row-sums controlled, we relate $\widehat{A}$ and $A$ exactly through a decomposition that isolates the contributions of $\Delta$ in both the numerator and denominator.
Lemma 10 (Exact perturbation identity and bound).
$\widehat{A}-A\;=\;\widehat{D}^{-1}\Delta\;+\;(\widehat{D}^{-1}-D^{-1})S.$
Moreover, whenever $\|E\|_{2}<s_{\min}$,
$\|\widehat{A}-A\|_{2}\;\leq\;\frac{\|\Delta\|_{2}}{s_{\min}-\|E\|_{2}}\;+\;\frac{\|S\|_{2}\,\|E\|_{2}}{s_{\min}(s_{\min}-\|E\|_{2})}.$
Proof.
Using $\widehat{S}=S+\Delta$ and $\widehat{D}=D+E$,
$\widehat{A}-A=\widehat{D}^{-1}\widehat{S}-D^{-1}S=\widehat{D}^{-1}\Delta+(\widehat{D}^{-1}-D^{-1})S.$
For the bound, we apply the submultiplicativity property of norms. When $\|E\|_{2}<s_{\min}$, we have $\|D^{-1}\|_{2}=1/s_{\min}$ and $\|\widehat{D}^{-1}\|_{2}\leq 1/(s_{\min}-\|E\|_{2})$. Moreover,
$\|\widehat{D}^{-1}-D^{-1}\|_{2}=\|\widehat{D}^{-1}(D-\widehat{D})D^{-1}\|_{2}\leq\|\widehat{D}^{-1}\|_{2}\,\|E\|_{2}\,\|D^{-1}\|_{2}\leq\frac{\|E\|_{2}}{s_{\min}(s_{\min}-\|E\|_{2})}.$
Hence,
$\|\widehat{A}-A\|_{2}\leq\|\widehat{D}^{-1}\|_{2}\,\|\Delta\|_{2}+\|\widehat{D}^{-1}-D^{-1}\|_{2}\,\|S\|_{2}\leq\frac{\|\Delta\|_{2}}{s_{\min}-\|E\|_{2}}+\frac{\|S\|_{2}\,\|E\|_{2}}{s_{\min}(s_{\min}-\|E\|_{2})}.$
Specializing Lemma 10 to the regime $\|E\|_{2}\leq s_{\min}/2$ (guaranteed w.h.p. by Lemma 9), we obtain a concise spectral bound for $\|\widehat{A}-A\|_{2}$ in terms of $\|\Delta\|_{2}$.
Lemma 11 (Attention deviation).
If $\|E\|_{2}\leq s_{\min}/2$, then
$\|\widehat{A}-\!A\|_{2}\leq\frac{2\|\Delta\|_{2}}{s_{\min}}+\frac{2\|S\|_{2}}{s_{\min}^{2}}\,\sqrt{N}\,\|\Delta\|_{2}.$
Proof.
From Lemma 10,
$\|\widehat{A}-A\|_{2}\leq\frac{\|\Delta\|_{2}}{s_{\min}-\|E\|_{2}}+\frac{\|S\|_{2}\,\|E\|_{2}}{s_{\min}(s_{\min}-\|E\|_{2})}.$
Since $\|E\|_{2}\leq s_{\min}/2$, it follows that
$\frac{1}{s_{\min}-\|E\|_{2}}\;\leq\;\frac{1}{s_{\min}/2}=\frac{2}{s_{\min}}.$
Substituting this bound gives
$\|\widehat{A}-A\|_{2}\;\leq\;\frac{2\|\Delta\|_{2}}{s_{\min}}+\frac{2\|S\|_{2}}{s_{\min}^{2}}\,\|E\|_{2}$
By Lemma 8(1), $\|E\|_{2}\leq\|\Delta\|_{\infty}\leq\sqrt{N}\,\|\Delta\|_{2}$. Therefore,
$\|\widehat{A}-A\|_{2}\;\leq\;\frac{2\|\Delta\|_{2}}{s_{\min}}+\frac{2\|S\|_{2}}{s_{\min}^{2}}\,\sqrt{N}\,\|\Delta\|_{2},$
which proves the claim. ∎
Finally, we translate attention deviation into end-to-end output deviation by a single multiplication with the value matrix $V$, yielding the main finite-sample guarantee.
Theorem 12 (End-to-end output error).
Let $V\in\mathbb{R}^{N\times d}$ be the value matrix. With probability at least $1-\delta$, if $\|E\|_{2}\leq s_{\min}/2$ then
$\|\widehat{O}-O\|_{F}\;\leq\;\Big(\tfrac{2}{s_{\min}}+\tfrac{2\|S\|_{2}\sqrt{N}}{s_{\min}^{2}}\Big)\left(\frac{4}{\sqrt{2\pi}}\,\frac{NP}{\beta}+\mathcal{O}\big(NPe^{-c\beta}\big)+4\,\frac{N}{\sqrt{L}}\,\sqrt{\log\!\tfrac{2N}{\delta}}+\tfrac{4}{3}\,\frac{N}{L}\,\log\!\tfrac{2N}{\delta}\right)\,\|V\|_{F},$
where $c=2\tanh(1)$.
Proof.
By the estimator identity, $\widehat{O}=\widehat{A}V$ and $O=AV$, hence using submultiplicativity of the Frobenius norm,
$\|\widehat{O}-O\|_{F}=\|(\widehat{A}-A)V\|_{F}\leq\|\widehat{A}-A\|_{2}\,\|V\|_{F}.$
Under the condition $\|E\|_{2}\leq s_{\min}/2$, Lemma 11 (Attention deviation) gives
$\|\widehat{A}-A\|_{2}\;\leq\;\Big(\tfrac{2}{s_{\min}}+\tfrac{2\|S\|_{2}\sqrt{N}}{s_{\min}^{2}}\Big)\,\|\widehat{S}-S\|_{2}.$
Applying Theorem 6 (Kernel deviation) yields, with probability at least $1-\delta$,
$\|\widehat{S}-S\|_{2}\;\leq\;\|\tilde{B}\|_{2}\;+\;4\,\frac{N}{\sqrt{L}}\,\sqrt{\log\!\frac{2N}{\delta}}\;+\;\frac{4}{3}\,\frac{N}{L}\,\log\!\frac{2N}{\delta}.$
Finally, substitute the explicit bias bound from Lemma 7:
$\|\tilde{B}\|_{2}\;\leq\;\frac{4}{\sqrt{2\pi}}\,\frac{NP}{\beta}\;+\;\mathcal{O}\big(NPe^{-c\beta}\big),\qquad c=2\tanh(1).$
Combining the three equations proves the stated inequality:
$\|\widehat{O}-O\|_{F}\;\leq\;\Big(\tfrac{2}{s_{\min}}+\tfrac{2\|S\|_{2}\sqrt{N}}{s_{\min}^{2}}\Big)\left(\frac{4}{\sqrt{2\pi}}\,\frac{NP}{\beta}+\mathcal{O}\big(NPe^{-c\beta}\big)+4\,\frac{N}{\sqrt{L}}\,\sqrt{\log\!\tfrac{2N}{\delta}}+\tfrac{4}{3}\,\frac{N}{L}\,\log\!\tfrac{2N}{\delta}\right)\,\|V\|_{F}.$ (6)
Proof of Theorem 2.
Under assumptions (A1) and (A2), we have $s_{\min}\geq C_{1}N$ and $\|S\|_{2}\leq C_{2}N$. Therefore, the prefactor on the r.h.s. of eq. (6) boils down to
$\frac{2}{s_{\min}}+\frac{2\|S\|_{2}\sqrt{N}}{s_{\min}^{2}}\;\leq\;\frac{2}{C_{1}N}+\frac{2C_{2}N\sqrt{N}}{C_{1}^{2}N^{2}}\;=\;\mathcal{O}\Big(\frac{1}{\sqrt{N}}\Big).$ (7)
Therefore, combining eqs. (6) and (7) yields
$\|\widehat{O}-O\|_{F}=$ $\mathcal{O}\Big(\tfrac{1}{\sqrt{N}}\Big)\!\Big(\tfrac{NP}{\beta}+\tfrac{N}{\sqrt{L}}\sqrt{\log\tfrac{2N}{\delta}}+\tfrac{N}{L}\log\tfrac{2N}{\delta}+NPe^{-c\beta}\Big)\|V\|_{F}$
$=$ $\mathcal{O}\Big(\tfrac{P\sqrt{N}}{\beta}+\sqrt{\tfrac{N}{L}\log\tfrac{2N}{\delta}}+\tfrac{\sqrt{N}}{L}\log\tfrac{2N}{\delta}+P\sqrt{N}\,e^{-c\beta}\Big)\|V\|_{F}.$ (8)
Dividing both sides of eq. (8) by $\sqrt{N}$ to express the bound in terms of the per-token RMS error gives
$\|\widehat{O}-O\|_{\mathrm{rms}}\;\leq\;\mathcal{O}\Big(\tfrac{P}{\beta}+\sqrt{\tfrac{\log(2N/\delta)}{L}}+\tfrac{1}{L}\log\tfrac{2N}{\delta}+Pe^{-c\beta}\Big)\|V\|_{F}.$
To compare the last two variance terms, observe that
$\frac{\tfrac{1}{L}\log\tfrac{2N}{\delta}}{\sqrt{\tfrac{\log(2N/\delta)}{L}}}\;=\;\sqrt{\tfrac{\log(2N/\delta)}{L}}.$
Hence, whenever $L\geq\log(2N/\delta)=\Theta(\log N)$, the $(1/L)\log(2N/\delta)$ term is asymptotically dominated by $\sqrt{\log(2N/\delta)/L}$ and can therefore be absorbed into it. The exponentially small correction $Pe^{-c\beta}$ is also negligible for moderate values of $\beta$. Absorbing all constants and the mild difference between $\log(2N/\delta)$ and $\log(N/\delta)$ into the Big-$\mathcal{O}$, we obtain
$\|\widehat{O}-O\|_{\mathrm{rms}}\;=\;\mathcal{O}\!\Big(\tfrac{P}{\beta}+\sqrt{\tfrac{\log(N/\delta)}{L}}\Big)\|V\|_{F},$
with probability at least $1-\delta$. This completes the proof. ∎
7 Causal Race Attention
Algorithm 2 RACE Attention (causal)
Input: $Q,K,V\in\mathbb{R}^{N\times d}$; number of hash tables $L$; number of hyperplanes $P$; temperature $\beta>0$.
Output: $\widehat{O}\in\mathbb{R}^{N\times d}$.
1:for $\ell=1,\dots,L$ do
2:  Draw $W^{(\ell)}\in\mathbb{R}^{P\times d}$ // $P$ random hyperplanes
3:  Define the corner set $\mathcal{V}=\{\pm 1\}^{P}$ ($R=2^{P}$) with $v_{r}\in\{\pm 1\}^{P}$ // $R$ corners
4:  Build $\Phi_{Q}^{(\ell)},\Phi_{K}^{(\ell)}\in\mathbb{R}^{N\times R}$ with rows
$\phi^{(\ell)}(x)]_{r}=\frac{\exp\{\beta\,(\tanh(W^{(\ell)}x))^{\top}v_{r}\}}{\sum_{r^{\prime}}\exp\{\beta\,(\tanh(W^{(\ell)}x))^{\top}v_{r^{\prime}}\}},\quad x\in\{Q_{i},K_{j}\}.$
5:  Initialize cumulative bucket statistics:
$A_{\mathrm{cum}}^{(\ell)}\leftarrow\mathbf{0}_{R}\in\mathbb{R}^{R},\qquad B_{\mathrm{cum}}^{(\ell)}\leftarrow\mathbf{0}_{R\times d}\in\mathbb{R}^{R\times d}.$
6:  for $t=1,\dots,N$ do
7:   $\Phi_{K}^{(\ell)}[t,:]\in\mathbb{R}^{R}$,  $V_{t}\in\mathbb{R}^{d}$
8:   $A_{\mathrm{cum}}^{(\ell)}\leftarrow A_{\mathrm{cum}}^{(\ell)}+(\Phi_{K}^{(\ell)}[t,:])^{\top}$ // $\mathbb{R}^{R}$
9:   $B_{\mathrm{cum}}^{(\ell)}\leftarrow B_{\mathrm{cum}}^{(\ell)}+(\Phi_{K}^{(\ell)}[t,:])^{\top}V_{t}$ // $\mathbb{R}^{R\times d}$
10:   $\Phi_{Q}^{(\ell)}[t,:]\in\mathbb{R}^{R}$
11:   $\texttt{num}_{t}^{(\ell)}\leftarrow\Phi_{Q}^{(\ell)}[t,:]\;B_{\mathrm{cum}}^{(\ell)}$ // $(1\times R)\cdot(R\times d)=\mathbb{R}^{d}$
12:   $\texttt{den}_{t}^{(\ell)}\leftarrow\Phi_{Q}^{(\ell)}[t,:]\;A_{\mathrm{cum}}^{(\ell)}$ // $(1\times R)\cdot(R)=\mathbb{R}$
13:  end for
14:end for
15:For each $t=1,\dots,N$:
$\texttt{Num}_{t}\;=\;\frac{1}{L}\sum_{\ell=1}^{L}\texttt{num}_{t}^{(\ell)}\in\mathbb{R}^{d},\qquad\texttt{Den}_{t}\;=\;\frac{1}{L}\sum_{\ell=1}^{L}\texttt{den}_{t}^{(\ell)}\in\mathbb{R},\qquad\widehat{O}_{t}\;=\;\frac{\texttt{Num}_{t}}{\texttt{Den}_{t}}\in\mathbb{R}^{d}.$
16:return $\widehat{O}=\begin{bmatrix}\widehat{O}_{1}^{\top}\\ \vdots\\ \widehat{O}_{N}^{\top}\end{bmatrix}\in\mathbb{R}^{N\times d}$.
We implemented the causal version efficiently using OpenMP-based parallelization rather than a naïve nested-loop approach. Each hash table is processed in a separate thread with its own cumulative bucket arrays, and updates are performed incrementally in a single left-to-right scan. This avoids redundant recomputation at every step using torch.cumsum() and enables CPU-level parallel execution with negligible synchronization overhead.
Figure 6: RACE Attention pipeline from the inputs $Q,K,V\in R^{N\times d}$ to the output $O\in R^{N\times d}$: queries/keys are soft-hashed into $R$ buckets across $L$ tables and $M$ ensembles, keys/values form per-bucket summaries, and each query mixes the matched summaries to produce $O$.
Figure 7: An intuitive schematic of how RACE Attention runs with $L$ hash tables and $R$ buckets per table. Similarity between Queries and Keys is highest if they both hash to same buckets across all hash tables.